Shifter leader was hy Shifters were 1. we had beam at the beginning of the shift, and have requested to dump the beam since last shift. Collision rate was very low, while until 11:30, the beam was dumped under the request from all experiments. Up to now, we have not got the beam back. And after the beam was dumped, the D5 field has been ramped up from 350 to 700. FS is 2.3@1723 2. data run: 16282(from last shift), 16283(very short one), 16284, 16285, 16288, 16289(short one before they dump the beam). pedestal run: TPC 16290, non-TPC pedestal run 16291 3. No new detector problem so far. _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Tue Jun 20 16:02:56 2006
This archive was generated by hypermail 2.1.8 : Tue Jun 20 2006 - 16:03:10 EDT